OpenStudy (anonymous):

A boat costs $16,600 and decreases in value by 14% per year. How much will the boat be worth after 11 years? (A)$2,717.00 (B)$70,155.46 (C)$3,159.30 (D)$16,446.00

OpenStudy (tkhunny):

Really Big Hint: 1 - 14% = 1 - 0.14 = 0.86

OpenStudy (anonymous):

i got A..did I do it right? @tkhunny

OpenStudy (tkhunny):

What was your answer and how did you get it? Start: 16600 Year 1: 16600*0.86 Year 2: 16600*0.86*0.86 = 16600 * 0.86^2

OpenStudy (anonymous):

I did what you did but i did it til i got to 11 years. @tkhunny

OpenStudy (anonymous):

i think i did something wrong..

OpenStudy (tkhunny):

Did you notice that you can skip straight to year 11? Year 11: 16600*0.86^11 = 16600*0.1903193578437 = 3159.30